OpenStudy (anonymous):

how do you integrate cos(x^1/2)dx

OpenStudy (mertsj):

Let u = sqrt(x)

OpenStudy (lalaly):

\[\int\limits{\cos(x^{\frac{1}{2}})dx}=\int\limits{\cos(\sqrt x)dx}\]let u= sqrt(x) .. du=1/(2sqrtx) so the integration becomes\[\int\limits{2ucos(u)du}\]now integrate by parts